The velocity field of a flow is given 2 2 V x ti y t x t j m s = + −+ 2 4 ( 1) 2 / ⎡ ⎤ ⎣ ⎦ K JK , where x and y are in meters and t is seconds. For fluid particles on the x axis, determine the speed and direction of flow.


1
Expert's answer
2021-10-11T07:01:40-0400

velocity field of a flow: In continuum mechanics the flow velocity in fluid dynamics, also macroscopic velocity

If a fluid particle moves along the curve

"x\n(\nt\n)\n=\n(\nx\n(\nt\n)\n,\ny\n(\nt\n)\n)"

then its velocity at time t  is the derivative

"\mathbf{v}= \frac{d\mathbf{x}}{dt}"

of its position with respect to t. Thus, for a time-independent velocity vector field

"\mathbf{v}(x, y) = ( v_1(x, y), v_2(x, y) )"

the fluid particles will move in accordance with an autonomous, first order system of ordinary differential equations



"\frac{dx}{dt}= v_1(x, y),\qquad \frac{dy}{dt}= v_2(x, y)."